# Automated Retinal Layer and Fluid Segmentation and Cross-sectional Analysis using Spectral Domain Optical Coherence Tomography Images for Diabetic Retinopathy

### v1

@@ II-E Training Settings
We employed 5-fold cross-validation, stratified by DR diagnosis, with a 4:1:1 ratio for training, validation, and testing. Each input consisted of a 3-channel image constructed by three repetitions of a single B-scan. Each training B-scan is resized to $512\times 512$. To mitigate class imbalance, excess Vitreous and Choroid regions were cropped. Various augmentation techniques are applied, including lateral flipping, Gaussian noise addition, contrast enhancement, rotation within the B-scan plane, and random intensity shifting.
@@
For loss functions, we use the combinations of Dice loss, cross-entropy (CE) loss, and the L1 loss of texture differences. Given the ground truth label $y$ and predicted label $\hat{y}$, for every pixel $i$, the Dice loss is calculated as:
@@
We empirically assign different class weights to CE loss to emphasize the class imbalance issue. Specifically, we assign 0.1 to Vitreous and Choroid, 0.5 to the rest of the layers, and 1 to fluid and HRF. We use AdamW optimizer with CosineAnnealing scheduler with the warm restart. We use the distributed parallel learning supported by the PyTorch Lightning module [^1] https://lightning.ai/, with a batch size of 8 and a learning rate of 1e-4. The training is deployed on NVIDIA V100 Volta GPU allocated by Cedar Compute Canada [^2] More information can be found at: https://docs.alliancecan.ca/wiki/Cedar.
II-F Evaluation Metrics
We evaluate the segmentation performance by overlapping areas and boundary alignment. We use the Dice similarity coefficient (DSC) to measure the similarity between the predicted and ground truth masks. Given correctly predicted pixels as True Positives(TP), incorrectly predicted pixels as False Positives(FP), and missing predicted pixels as False Negatives(FN), the Dice score is calculated as:
@@ II-F Evaluation Metrics
Additionally, we define the Under-Segmentation Score (USS) and the Over-Segmentation Score (OSS) to evaluate if the model fails to detect certain regions or assigns excessive labels to a class. Given the confusion matrix for N classes:
@@
We compute the USS and OSS for certain class C as:
@@
A higher USS score indicates that a significant portion of the ground truth class $C$ was not detected, leading to under-segmentation, while a higher OSS score suggests that the model over-predicts class $C$, leading to over-segmentation.
II-G Thickness Analysis
Layer thickness computation is performed using the K-Nearest Neighbors (K-NN) algorithm. The layer boundaries are converted into 3D point clouds. For each data point on the upper layer, the closest corresponding point on the lower layer is identified based on Euclidean distance. Due to their unbounded nature on one side, the Vitreous and Choroid layers are excluded from these calculations. Given the anatomical complexity of the foveal pit, the central region thickness is excluded from the analysis to ensure more reliable and interpretable measurements.
The thickness of fluid and hyperreflective foci (HRF) regions is computed through direct accumulation, which is subsequently projected onto an en-face view for visualization. The Early Treatment Diabetic Retinopathy Study (ETDRS) grid is employed to assess thickness variations systematically across different macular regions. As depicted in Figure 3, this grid divides the macula into three concentric circles with diameters of 1mmmm\$\mathrm{m}$m, 3mmmm\$\mathrm{m}$m, and 6mmmm\$\mathrm{m}$m, all centered on the fovea. These circles define the central, inner, and outer subfields, subdivided into four quadrants: superior, inferior, nasal, and temporal.
Figure 3: ETDRS diagram for both left and right eyes. The diameters of the central circle, inner ring, and outer ring were 1 mm, 3 mm, and 6 mm, respectively. Nine subfields are numbered and named as follows: 1-CS(Central Subfield), 2-SI(Superior Inner), 3-NI(Nasal Inner), 4-II(Inferior Inner), 5-TI(Temporal Inner), 6-SO(Superior Outer), 7-NO(Nasal Outer), 8-IO(Inferior Outer), 9-TO(Temporal Outer).
@@ III-A Segmentation
Table II presents the segmentation results for four models, with values averaged across five-fold cross-validation. Tables II(a) and II(b) separately report the DSC and NSD metrics. A repeated-measures analysis of variance (ANOVA) is conducted to compute the F-statistics and p-values for each retinal layer in both DSC and NSD. The ”Avg” column represents the average performance across all layers, with the highest score for each layer highlighted in bold.
SwinUNETR achieves the highest overall DSC and NSD among the evaluated models, demonstrating superior segmentation performance, particularly in the OPL, Choroid, and HRF regions. VM-UNet exhibits competitive performance, achieving the best DSC and NSD scores in the Vitreous, RNFL, and fluid regions. U-Net and SegFormer perform comparably, though U-Net slightly outperforms SegFormer in DSC across most layers, whereas SegFormer demonstrates marginally better NSD performance.
Notably, high F-values and small p-values are observed in the RNFL, GCL+IPL, INL, ONL+IS, fluid, and HRF layers, indicating significant variability in segmentation accuracy across models within these regions. These findings suggest that model predictions are less consistent in these layers, potentially due to structural complexity or segmentation challenges inherent to these regions.
Model Layer Dice Vitreous RNFL GCL+IPL INL OPL ONL+IS EZ OS RPE Choroid Fluid HRF Avg.
@@
F-value 18.86 161.84 95.87 50.10 37.33 58.13 14.44 13.58 13.74 5.569 33.97 52.64 -
p-value 4.130×$\times 10^{-5}$ 2.504×$\times 10^{-10}$ 5.322×$\times 10^{-5}$ 2.166×$\times 10^{-7}$ 1.110×$\times 10^{-6}$ 9.368×$\times 10^{-8}$ 1.545×$\times 10^{-4}$ 2.072×$\times 10^{-4}$ 1.961×$\times 10^{-4}$ 9.023×$\times 10^{-3}$ 1.861×$\times 10^{-6}$ 1.640×$\times 10^{-7}$
(a) Mean Dice Similarity Coefficient (DSC) with repeated-measure ANOVA results
Model Layer NSD Vitreous RNFL GCL+IPL INL OPL ONL+IS EZ OS RPE Choroid Fluid HRF Avg.
@@
F-value 27.65 36.34 31.55 22.64 11.30 4.054 19.06 12.36 3.485 6.568 19.62 44.15 -
p-value 5.640×$\times 10^{-6}$ 1.288×$\times 10^{-6}$ 2.776×$\times 10^{-6}$ 1.614×$\times 10^{-5}$ 4.901×$\times 10^{-4}$ 2.632×$\times 10^{-2}$ 3.915×$\times 10^{-5}$ 3.238×$\times 10^{-4}$ 4.142×$\times 10^{-2}$ 4.864×$\times 10^{-3}$ 3.372×$\times 10^{-5}$ 4.389×$\times 10^{-7}$
(b) Mean Normalized Surface Dice (NSD) with repeated-measure ANOVA results
@@
TABLE II: Comparison of segmentation performance across four models. Dice and NSD scores are calculated by averaging over five validation folds. The ”Average” column represents the mean performance across all regions per model. The best score in each layer is highlighted in bold. The bottom rows show the F-value and p-value from repeated-measures ANOVA for statistical significance.
Figure 4 presents the mean difference in DSC and NSD between each model pair across retinal regions. For a model pair denoted as ”A vs B,” the mean difference is computed by subtracting model A’s mean performance from model B’s. The significance of these differences is indicated by an asterisk marker preceding the model pair: ”***” denotes strong statistical significance ($p<0.001$), ”**” represents moderate significance ($p<0.01$), and ”*” indicates statistical significance at $p<0.05$. The absence of a marker signifies no significant difference.
Significant mean differences with strong statistical significance are observed in the fluid and HRF regions. SwinUNETR and VM-UNet consistently demonstrate superior performance across retinal layers. While U-Net and SegFormer perform competitively in most layers, their performance in fluid and HRF regions is inferior to that of SwinUNETR and VM-UNet.
SwinUNETR and VM-UNet exhibit comparable performance in the INL, OPL, Choroid, and fluid regions for DSC. However, VM-UNet significantly outperforms SwinUNETR in the Vitreous, RNFL, GCL+IPL, and ONL+IS layers, whereas SwinUNETR achieves considerably better performance than VM-UNet in the EZ, OS, RPE, and HRF regions. Regarding NSD, SwinUNETR demonstrates a significantly larger NSD than VM-UNet in the Vitreous, OPL, ONL+IS, RPE, Choroid, and HRF layers, with no significant differences observed in the remaining regions.
Figure 4: Comparison of the mean difference of DSC and NSD scores between models for each retinal region. For each pair of models shown in the subplot, the mean difference is calculated by subtracting the mean of the first model from the second model. The marker before the name of each model pair indicates different significant levels(p-values) calculated by the Tukey HSD method. ***: $p<0.001$, **: $p<0.01$, *: $p<0.05$, no marker: $p\geq 0.05$. Some barplots are invisible due to the limited precision of the mean difference.
@@
Figure 5 presents examples of segmentation model predictions. Each sub-image displays five B-scans selected from the 60 central B-scans. The predicted segmentation is generated using the sub-fold model corresponding to the test set to which the patient belongs.
@@
Figure 5(a) illustrates an NPDR patient exhibiting severe intraretinal fluid. All four models successfully segment the majority of fluid regions. However, U-Net and SegFormer demonstrate weaker fine-layer segmentation performance than SwinUNETR and VM-UNet, particularly in the RNFL and OPL layers. VM-UNet excels in preserving layer continuity and structural integrity, whereas U-Net and SwinUNETR exhibit discontinuities in the OPL layer in the fourth and fifth B-scans. Notably, VM-UNet tends to under-segment the fluid relative to the other models.
@@
Figure 5(b) displays an NPDR patient with pronounced HRF. Shading artifacts beneath large HRF clusters disrupt the continuity of the lower layers. VM-UNet demonstrates superior performance in maintaining layer integrity despite losing pixel intensity in the bottom three B-scans. In contrast, U-Net and SegFormer struggle to compensate for these artifacts, while SwinUNETR erroneously misclassifies portions of HRF within the Choroidal region.
@@
Figure 5(c) depicts a PDR patient with severe fluid accumulation. U-Net exhibits the weakest performance in fluid segmentation among all models, particularly in the third and fourth B-scans. Additionally, all models show varying degrees of under-segmentation in the final B-scan.
Figure 5(d) illustrates a PDR patient with partial retinal detachment. VM-UNet demonstrates robust performance across all Example B-scans, providing a more accurate segmentation of the RNFL compared to the other models.
(a) NPDR patient with severe fluid accumulation.
@@
(d) PDR patient with partial retinal detachment.
@@
Figure 6 shows the USS and OSS for the top-2 performance models SwinUNETR and VM-UNet. VM-UNet shows lower USS than SwinUNETR in most regions except for HRF, and it has less over-segmentation in most areas except for INL, OPL, EZ and OS. Overall, both models tend to under-segment in OPL, EZ and OS layers, plus the fluid and HRF. Over-segmentation is observed in the INL, OPL, ONL+IS, EZ and OS layers.
@@ III-B Thickness
Statistical significance between the retinal thickness and DR groups is analyzed in Table III. The thickness analysis is applied to the predicted segmentation of the top-2 performance models SwinUNETR and VM-UNet for comparison with the ground-truth segmentation. For each layer sector, the p-values of Welch’s t-test (first row) and Analysis of Covariance (ANCOVA) adjusted by age, gender and diabetes duration (second row) are calculated. The p-values less than 0.05 are considered statistically significant. For ground-truth results, PDR has a significantly larger thickness than NPDR in the NI and II sectors of RNFL before and after p-value adjustment, but vice versa in II of GCL+IPL and all sectors of the OS layer except TI. NPDR has a significantly larger thickness in TI and IO of the GCL+IPL layer, TI of the OS layer, and more extensive volume accumulation in NI and TI of HRF. However, the significance is lost after adjusting for confounders. No significance is found in other layer sectors.
For model prediction results, SwinUNETR shows no significance in RNFL, while VM-Unet preserves the significance of ground-truth results but exaggerates them in the SI, SO, and TO sectors. SwinUNETR shows consistent significance in the IO sector of GCL+IPL, while VM-Unet only shows significance in II. Both models maintain the significance of the OS layer in most sectors except for NI and NO sectors but hyperbolize in the IO sector of RPE. No significance is shown in the fluid and HRF region for both models.
@@
Fig 7 shows the thickness distribution across different retinal layers and sectors for NPDR and PDR groups. The outliers are removed outside the $5^{th}$ and $95^{th}$ percentiles, allowing a more robust interpretation of group differences. The diamond markers in each subplot show the mean thickness of each DR group without outlier removal. The PDR group has larger mean values and broader distributions in most sectors of RNFL, INL, OPL, EZ, and fluid, but it is vice versa for GCL+IPL, ONL+IS, OS, RPE and HRF.
Region Sector
SI NI II TI SO NO IO TO
Ground-truth
RNFL 0.152 0.045 0.110 0.011 0.094 0.053 0.448 0.537
0.183 0.033 0.083 0.012 0.133 0.052 0.550 0.623
GCL+IPL 0.201 0.364 0.025 0.050 0.865 0.536 0.045 0.246
0.230 0.447 0.035 0.084 0.764 0.604 0.060 0.377
INL 0.943 0.886 0.255 0.117 0.935 0.628 0.283 0.634
0.906 0.714 0.276 0.069 0.990 0.676 0.274 0.514
OPL 0.663 0.263 0.525 0.693 0.331 0.875 0.647 0.102
0.833 0.288 0.511 0.744 0.325 0.867 0.710 0.089
ONL+IS 0.314 0.713 0.101 0.218 0.109 0.910 0.105 0.630
0.333 0.588 0.072 0.228 0.101 0.724 0.070 0.674
EZ 0.722 0.759 0.562 0.884 0.921 0.498 0.737 0.468
0.649 0.652 0.503 0.702 0.998 0.496 0.807 0.500
OS 0.001 0.010 0.003 0.042 0.000 0.001 0.000 0.001
0.003 0.016 0.006 0.062 0.000 0.002 0.000 0.002
RPE 0.517 0.136 0.252 0.168 0.428 0.212 0.284 0.996
0.498 0.828 0.293 0.740 0.518 0.809 0.252 0.568
Fluid 0.517 0.136 0.252 0.168 0.428 0.212 0.284 0.996
0.517 0.101 0.213 0.159 0.519 0.283 0.381 0.825
HRF 0.535 0.634 0.033 0.250 0.032 0.114 0.322 0.572
0.681 0.528 0.059 0.284 0.072 0.120 0.384 0.580
SwinUNETR / VM-UNet
RNFL 0.183 / 0.039 0.094 / 0.027 0.349 / 0.142 0.504 / 0.017 0.076 / 0.028 0.149 / 0.089 0.594 / 0.777 0.059 / 0.007
0.201 / 0.039 0.063 / 0.014 0.259 / 0.107 0.518 / 0.020 0.110 / 0.042 0.125 / 0.076 0.607 / 0.861 0.062 / 0.007
GCL+IPL 0.387 / 0.436 0.461 / 0.408 0.053 / 0.018 0.135 / 0.158 0.921 / 0.726 0.413 / 0.296 0.044 / 0.068 0.463 / 0.435
0.401 / 0.478 0.543 / 0.507 0.075 / 0.024 0.213 / 0.263 0.918 / 0.679 0.397 / 0.240 0.048 / 0.080 0.603 / 0.611
INL 0.396 / 0.612 0.739 / 0.813 0.505 / 0.238 0.684 / 0.124 0.657 / 0.857 0.915 / 0.938 0.224 / 0.460 0.424 / 0.213
0.419 / 0.732 0.853 / 0.607 0.546 / 0.208 0.613 / 0.057 0.590 / 0.841 0.868 / 0.908 0.204 / 0.453 0.357 / 0.156
OPL 0.453 / 0.859 0.222 / 0.397 0.359 / 0.719 0.242 / 0.696 0.983 / 0.736 0.947 / 0.855 0.204 / 0.271 0.422 / 0.326
0.494 / 0.996 0.201 / 0.421 0.352 / 0.701 0.224 / 0.777 0.967 / 0.792 0.929 / 0.868 0.212 / 0.321 0.362 / 0.249
ONL+IS 0.154 / 0.468 0.729 / 0.937 0.117 / 0.364 0.087 / 0.128 0.083 / 0.150 0.712 / 0.982 0.023 / 0.066 0.427 / 0.497
0.166 / 0.495 0.564 / 0.967 0.086 / 0.298 0.110 / 0.155 0.061 / 0.120 0.462 / 0.838 0.014 / 0.053 0.482 / 0.578
EZ 0.329 / 0.339 0.564 / 0.466 0.330 / 0.153 0.179 / 0.230 0.558 / 0.216 0.279 / 0.681 0.549 / 0.312 0.668 / 0.908
0.474 / 0.464 0.667 / 0.622 0.409 / 0.252 0.368 / 0.437 0.613 / 0.231 0.229 / 0.629 0.538 / 0.344 0.649 / 0.974
OS 0.003 / 0.003 0.063 / 0.227 0.024 / 0.030 0.026 / 0.029 0.000 / 0.000 0.342 / 0.229 0.000 / 0.000 0.001 / 0.000
0.004 / 0.005 0.057 / 0.203 0.029 / 0.040 0.031 / 0.034 0.000 / 0.001 0.272 / 0.185 0.000 / 0.000 0.002 / 0.001
RPE 0.347 / 0.100 0.279 / 0.112 0.344 / 0.111 0.539 / 0.256 0.350 / 0.205 0.313 / 0.411 0.032 / 0.018 0.169 / 0.286
0.374 / 0.110 0.308 / 0.102 0.400 / 0.129 0.536 / 0.342 0.319 / 0.206 0.259 / 0.349 0.033 / 0.020 0.149 / 0.354
Fluid 0.277 / 0.599 0.076 / 0.108 0.183 / 0.353 0.108 / 0.158 0.916 / 0.128 0.910 / 0.384 0.125 / 0.497 0.486 / 0.416
0.277 / 0.618 0.053 / 0.083 0.159 / 0.322 0.100 / 0.154 0.864 / 0.165 0.983 / 0.418 0.067 / 0.532 0.386 / 0.345
HRF 0.716 / 0.845 0.564 / 0.681 0.061 / 0.125 0.523 / 0.694 0.064 / 0.068 0.353 / 0.113 0.467 / 0.279 0.871 / 0.788
0.880 / 0.973 0.432 / 0.560 0.102 / 0.171 0.524 / 0.714 0.120 / 0.129 0.343 / 0.101 0.553 / 0.367 0.835 / 0.776
TABLE III: Statistical significance of retinal thickness plus fluid and HRF volume differences for NPDR vs PDR. The first table shows results derived from the ground-truth label. The second table shows the results using the predicted segmentation, where each cell of two p-values represents the results of SwinUNETR and VM-UNet, respectively. Within each layer, the value on the first row represents the p-value from Welch’s t-test, and the second row represents the p-values adjusted by age, gender and duration of diabetes using Analysis of Covariance (ANCOVA). The p-values less than 0.05 are highlighted in bold.
Figure 9 presents four examples of thickness comparisons between NPDR and PDR groups using the ETDRS diagram described in Figure 3. Figures 8(a)9(b) correspond to cases of a 54-year-old male OD, a 57-year-old male OS, and a 59-year-old female OD and OS images, respectively. Each pair of patients is matched by age, gender, and eye laterality.
For each retinal region, the first row displays the En Face image overlaid with the corresponding layer thickness heatmap, while the second row presents the sector-wise quantitative average thickness. The En Face image is generated using each layer’s maximum intensity projection (MIP). For fluid and HRF, the En Face projection is derived from the entire retinal body (from the ILM to the BM), with thickness representing the accumulated volume in $\mu m^{3}$.
In the Example shown in Figure 8(a), PDR exhibits greater thickness than NPDR in nearly all RNFL sectors, except for the inner nasal and inferior regions. Similar trends are observed in the nasal and inferior OPL regions, inner sectors, and outer nasal sectors of fluid. Conversely, NPDR shows greater thickness in the inner sectors of the GCL+IPL, the inner superior INL, all OS sectors, and the temporal and superior HRF sectors.
In Figure 8(b), PDR demonstrates increased thickness in the nasal and superior RNFL sectors, the inner INL and OPL regions, nearly all fluid regions (except for the outer superior and temporal sectors), and all HRF regions. In contrast, NPDR exhibits greater thickness in the inner temporal and superior GCL+IPL regions and across all OS sectors.
For female patients shown in Figures 9(a) and 9(b), RNFL thinning is observed in the outer nasal and temporal sectors for PDR compared to NPDR. Additionally, GCL+IPL and RPE thinning are evident in the inner sectors of PDR. Fluid accumulation is significantly more pronounced in the inner sectors of PDR patients, whereas HRF is more prevalent in NPDR patients.
Figure 6: Comparison of USS(left) and OSS(right) for SwinUNETR and VM-UNet. Lower values indicate better segmentation performance.

### v3

@@ II-E Segmentation Model Training
We employed 5-fold cross-validation, stratified by DR diagnosis, with a 4:1:1 ratio for training, validation, and testing. Each input consisted of a 3-channel image constructed by three repetitions of a single B-scan. Each training B-scan was resized to $512\times 512$. To mitigate class imbalance among segmentation labels, excess Vitreous and Choroid regions were cropped. Various augmentation techniques were applied, including lateral flipping, Gaussian noise addition, contrast enhancement, rotation within the B-scan plane, and random intensity shifting.
@@
For loss functions, we used the combinations of Dice loss, cross-entropy (CE) loss, and the L1 loss of texture differences. Given the ground truth label $y$ and predicted label $\hat{y}$, for every pixel $i$, the Dice loss is calculated as:
@@
We empirically assigned different class weights to CE loss to emphasize the class imbalance issue. Specifically, we assigned 0.1 to Vitreous and Choroid, 0.5 to the rest of the layers, and 1 to fluid and HRF. We used AdamW optimizer with CosineAnnealing scheduler with the warm restart. We adopted the distributed parallel learning supported by the PyTorch Lightning module [^1] https://lightning.ai/, with a batch size of 8 and a learning rate of 1e-4. The training was deployed on NVIDIA V100 Volta GPU allocated by Cedar Compute Canada [^2] More information can be found at: https://docs.alliancecan.ca/wiki/Cedar.
II-F Semgmentation Performance Evaluation
We evaluated the segmentation performance by overlapping areas and boundary alignment. We used the Dice similarity coefficient (DSC) to measure the similarity between the predicted and ground truth masks. Given correctly predicted pixels as True Positives(TP), incorrectly predicted pixels as False Positives(FP), and missing predicted pixels as False Negatives(FN), the Dice score is calculated as:
@@ II-F Semgmentation Performance Evaluation
Additionally, we defined the Under-Segmentation Score (USS) and the Over-Segmentation Score (OSS) to evaluate if the model fails to detect certain regions or assigns excessive labels to a class. Given the confusion matrix for N classes:
@@
We computed the USS and OSS for certain class C as:
@@
A higher USS score indicates that a significant portion of the ground truth class $C$ is not detected, leading to under-segmentation, while a higher OSS score suggests that the model over-predicts class $C$, leading to over-segmentation. We used a heuristic cutoff value of 0.2 to determine if there is under-segmentation or over-segmentation.
II-G Retinal Layer Thickness Analysis
Layer thickness computation was performed using the K-Nearest Neighbors (K-NN) algorithm. The layer boundaries were converted into 3D point clouds. For each data point on the upper layer, the closest corresponding point on the lower layer was identified based on Euclidean distance. The distance is properly adjusted by the voxel dimension along each axis. The thickness maps are resized to the resolution of $350\times 350$ for consistent representation. The vitreous and choroid layers were excluded from these calculations due to their unbounded nature on one side. Given the anatomical complexity of the foveal pit, the central region thickness was excluded from the analysis to ensure more reliable and interpretable measurements.
The Early Treatment Diabetic Retinopathy Study (ETDRS) grid was employed to assess thickness variations systematically across different macular regions. As depicted in Figure 2, this grid divides the macula into three concentric circles with diameters of 1mmmm\$\mathrm{m}$m, 3mmmm\$\mathrm{m}$m, and 6mmmm\$\mathrm{m}$m, all centered on the fovea. These circles define the central, inner, and outer subfields, subdivided into four quadrants: superior, inferior, nasal, and temporal.
Figure 2: ETDRS diagram for both left and right eyes. The diameters of the central circle, inner ring, and outer ring are 1 mm, 3 mm, and 6 mm, respectively. Nine subfields are numbered and named as follows: 1-CS(Central Subfield), 2-SI(Superior Inner), 3-NI(Nasal Inner), 4-II(Inferior Inner), 5-TI(Temporal Inner), 6-SO(Superior Outer), 7-NO(Nasal Outer), 8-IO(Inferior Outer), 9-TO(Temporal Outer).
II-H Statistical Analysis
The DSC and NSD scores were calculated for model segmentation performance, and the thickness measurements were derived from the predicted segmentation. The mean DSC and NSD scores were compared within each retinal region. We used a generalized linear model (GLM) to assess the segmentation performance, thickness difference between the NPDR and PDR groups, and correlation between the thickness and visual acuity within each DR group.
Specifically, we compared the DSC and NSD scores of each pair of models with Gaussian distributions, and the performance is ranked via the effect size and p-values after the false discovery rate (FDR) correction. We compared the thickness differences between NPDR and PDR groups in each layer sector while adjusting for relevant covariates including age, gender, and duration of diabetes. Visual acuity was not included as it is considered a downstream clinical outcome rather than a demographic or biological confounder. The compound Poisson-Gamma distribution was used to model zero-inflated and highly skewed thickness measurements across DR groups while controlling for age, gender, and duration of diabetes. The correlation between the visual acuity and layer sector thickness within each DR group was modelled using Gaussian distribution while controlling for age, gender, and duration of diabetes. We converted each categorical variable to numerical values. We assigned 0 to NPDR and 1 to the PDR group, and assigned 0 to female and 1 to male. The models’ estimated coefficients (beta values) along with their 95% confidence intervals (CI) were calculated and visualized. Statistically significant results before and after FDR correction were explicitly highlighted.
@@ III-A Segmentation
Figure 3 illustrates a representative SD-OCT B-scan with ground truth retinal layer and fluid segmentation derived from the active-learning-based HITL semi-automatic segmentation pipeline. The segmentation delineates nine essential retinal layers: the Retinal Nerve Fiber Layer (RNFL), Ganglion Cell Layer and Inner Plexiform Layer (GCL+IPL), Inner Nuclear Layer (INL), Outer Plexiform Layer (OPL), Outer Nuclear Layer and Inner Segment Layer (ONL+IS), Ellipsoid Zone (EZ), Outer Segment Layer (OS), and Retinal Pigment Epithelium (RPE). The region above the Internal Limiting Membrane (ILM) is also identified as the Vitreous, while the Choroid lies beneath Bruch’s Membrane (BM). Fluid segmentation involves three primary fluid types: intraretinal fluid (IRF), subretinal fluid (SRF), and pigment epithelial detachment (PED), all of which appear as hypo-reflective regions between the ILM and BM. Furthermore, hyperreflective foci (HRF), which manifest as high-intensity dot-like or clustered lesions, are also segmented.
Figure 3: Example of retinal OCT and ground truth segmentation derived through the active-learning pipeline (image on the right) for a single B-scan(image on the left). This image was acquired from a 59-year-old female patient with NPDR. Ten retinal layers were segmented from top to bottom, plus the fluid and HRF within the retinal body.
Table II presents the segmentation results for four models, with values averaged across five-fold cross-validation. Tables II(a) and II(b) separately report the DSC and NSD metrics to quantify segmentation volume overlap and boundary distance respectively. SwinUNETR achieves the highest overall DSC and NSD among the evaluated models, demonstrating superior segmentation performance, particularly in the OPL, Choroid, and HRF regions. VM-UNet exhibits competitive performance, achieving the best DSC and NSD scores in the Vitreous, RNFL, and fluid regions. U-Net and SegFormer perform comparably, though U-Net slightly outperforms SegFormer in DSC across most layers, whereas SegFormer demonstrates marginally better NSD performance. These findings suggest that model predictions are less consistent in these layers, potentially due to structural complexity or segmentation challenges inherent to these regions.
Model Label Dice Vitreous RNFL GCL+IPL INL OPL ONL+IS EZ OS RPE Choroid Fluid HRF Avg.
@@
(a) Mean Dice Similarity Coefficient (DSC)
Model Label NSD Vitreous RNFL GCL+IPL INL OPL ONL+IS EZ OS RPE Choroid Fluid HRF Avg.
@@
(b) Mean Normalized Surface Dice (NSD)
@@
TABLE II: Comparison of segmentation performance across four models. Dice and NSD scores were calculated by averaging over five validation folds. The ”Average” column represents the mean performance across all retinal regions per model. The best score in each layer is highlighted in bold.
Figure 4 presents a comparative analysis of segmentation performance across U-Net, SegFormer, SwinUNETR, and VM-UNet using DSC (top row) and NSD (bottom row) across various retinal regions, excluding Vitreous and Choroid. Each boxplot illustrates the distribution of Dice metric per model, while the statistical significance of pairwise differences is determined using GLM with FDR correction. Significance markers below each region indicate statistically superior performance relative to other models: circles (‘o’) for U-Net, crosses (‘×’) for SegFormer, plus signs (‘+’) for SwinUNETR, and asterisks (‘*’) for VM-UNet. SwinUNETR and VM-UNet demonstrate consistent improvements over baseline models in several regions, notably in DSC and NSD for GCL+IPL, INL, Fluid, and HRF. However, U-Net and SegFormer show significantly better performance in the DSC of the EZ layer and NSD of the OS layer compared to SwinUNETR and VM-UNet. Additionally, SwinUNETR significantly outperforms VM-UNet in DSC for EZ, OS, RPE, and HRF, as well as NSD for OPL, ONL+IS, RPE, and HRF. Conversely, VM-UNet shows significantly better performance than SwinUNETR in DSC for RNFL and GCL+IPL. Quantitative results are shown in Supplementary Tables III and IV.
Figure 4: Comparison of segmentation performance across models using Area Dice (top) and Surface Dice (bottom) metrics for each retinal region (Vitreous and Choroid are excluded). Each boxplot summarizes the Dice scores for U-Net, SegFormer, SwinUNETR, and VM-UNet across all subjects. Statistical significance between models is assessed using GLM with FDR correction. Significance markers below each group indicate which models significantly outperform others: ’o’ indicates significantly better than U-Net, ’×’ better than SegFormer, ’+’ better than SwinUNETR, and ’*’ better than VM-UNet.
@@
Figure 5 presents some representative examples of segmentation model predictions. Each sub-image displays five B-scans selected from the 60 central B-scans. The predicted segmentation was generated using the sub-fold model corresponding to the test set to which the patient belongs.
@@
Figure 5(a) illustrates a representative NPDR patient exhibiting severe intraretinal fluid. All four models successfully segment the majority of fluid regions. However, U-Net and SegFormer demonstrate weaker fine-layer segmentation performance than SwinUNETR and VM-UNet, particularly in the RNFL and OPL layers. VM-UNet excels in preserving layer continuity and structural integrity, whereas U-Net and SwinUNETR exhibit discontinuities in the OPL layer in the fourth and fifth B-scans.
@@
Figure 5(b) displays a representative NPDR patient with pronounced HRF. Shading artifacts beneath large HRF clusters disrupt the continuity of the lower layers. VM-UNet demonstrates superior performance in maintaining layer integrity despite losing pixel intensity in the bottom three B-scans. In contrast, U-Net and SegFormer struggle to compensate for these artifacts, while SwinUNETR erroneously misclassifies portions of HRF within the Choroidal region. Notably, VM-UNet tends to under-segment the HRF relative to the other models.
@@
Figure 5(c) depicts a representative PDR patient with severe fluid accumulation. U-Net exhibits the weakest performance in fluid segmentation among all models, particularly in the third and fourth B-scans. Additionally, all models show varying degrees of under-segmentation in the final B-scan.
Figure 5(d) illustrates a representative NPDR patient with SRF. SwinUNETR and VM-UNet show superior SRF segmentation and surrounding layer boundary refinements than U-Net and SegFormer. SwinUNETR tends to over-segment both layers and fluid than VM-UNet in both layers and fluid, which is explicitly shown in the OPL layer and SRF of the fourth row.
(a) NPDR patient with severe fluid accumulation
@@
(d) NPDR patient with subretinal fluid (SRF).
@@
Figure 6 shows the USS and OSS for the top-2 performance models SwinUNETR and VM-UNet. VM-UNet shows lower USS than SwinUNETR in most regions except for HRF, and it has less over-segmentation in most areas except for INL, OPL, EZ and OS. Overall, using the 0.2 cutoff value, both models tend to under-segment in OPL, EZ and OS layers, plus the fluid and HRF. Over-segmentation is observed in the INL, OPL, ONL+IS, EZ and OS layers.
Figure 6: Comparison of USS(left) and OSS(right) for SwinUNETR and VM-UNet. Lower values indicate better segmentation performance. A heuristic cutoff value of 0.2 was used to determine if there are under-segmentations or over-segmentations.
@@ III-B Thickness
Fig 7 shows the violin plot about the distribution of the ground-truth-derived-thickness across different retinal layers and sectors for NPDR and PDR groups. The outliers are removed outside the $5^{th}$ and $95^{th}$ percentiles, allowing a more robust interpretation of group differences. The diamond markers in each subplot show the mean thickness of each DR group without outlier removal. The PDR group has larger mean values and broader distributions in most sectors of fluid and HRF, whereas the NPDR group has larger mean thickness in most sectors of GCL+IPL and all sectors of ONL+IS.